OpenStudy (athenaalien):

Derive this identity from the sum and difference formulas for cosine: cos a cos b=(1/2)[cos(a-b)+cos(a+b)] Please show the steps to solve this

satellite73 (satellite73):

do you know the formula for the sum?

satellite73 (satellite73):

\[\overbrace{\cos(a-b)}^{\text{put the formula here}}+\overbrace{\cos(a+b)}^{\text{and formula here}}\] add. and you will get it

OpenStudy (athenaalien):

What about this one: Use the trigonometric subtraction formula for sine to verify this identity: cos((pi/2)-x)=sinX @satellite73

satellite73 (satellite73):

same thing, use the subtraction angle formula for cosine \[\cos(a-b)=\cos(a)\cos(b)+\sin(a)\sin(b)\] put \(a=\frac{\pi}{2},b=x\) and compute

OpenStudy (athenaalien):

Oh ok a couple more questions: 1. If cosx=1/3 and x is in quadrant II, answer these: a. sin(x/2) b. tan(x/2) 2. Solve these equations graphically on the interval [0,pi]. Sketch the graph and list the solutions. a. sinx+1=cosx b. cos(2x)-1=sinx How would I solve these? @satellite73

satellite73 (satellite73):

ack who knows?

satellite73 (satellite73):

you need the half angle formula for sine for the first one

satellite73 (satellite73):

since you are in quadrant II, half of that angle is in quadrant 1, so sine is positive \[\sin(\frac{x}{2})=\sqrt{\frac{1-\cos(x)}{2}}\] replace \(\cos(x)\) by \(\frac{1}{3}\)

OpenStudy (athenaalien):

ok that makes sense

OpenStudy (athenaalien):

what would the equation for part b be? @satellite73

satellite73 (satellite73):

i don't know the one for tangent, you should google it if it isn't in the book can't memorize every damned thing!